# A Geometric Unification of Generative AI with Manifold-Probabilistic Projection Models

### v1

@@ 3 Geometric View of Diffusion Models
Since the domain of both the encoder $F$ and the distance $\mathcal{D}_{\mathcal{M}}$ is the ambient space $\mathbb{R}^{D}$, effectively training mappings that enable the diffusion-like flow from corrupted images back to clean ones on the manifold requires sampling the high-dimensional ambient space, which is an inherently challenging task due to the curse of dimensionality. Following the approach of diffusion models, we generate ambient samples by adding Gaussian noise to the data points. While this sampling strategy does not cover all possible corruptions, it empirically produces useful mappings. Notably, although the models are trained using Gaussian noise, they generalize well to other types of image corruption during testing.
@@
To connect a corrupted image to its clean projection, we use the concept of the score. The score is a $D$-dimensional vector field defined by $s(x)=\nabla_{x}\log P(x)$, which points in the direction of the steepest ascent of the probability density. For the distance-based probability distribution $P_{d}(x)$ defined in Eq. 2, we obtain:
@@
Since $\mathcal{D}_{\mathcal{M}}(x)$ is the distance to the manifold, its normalized gradient $\nabla_{x}\mathcal{D}_{\mathcal{M}}(x)/|\nabla_{x}\mathcal{D}_{\mathcal{M}}(x)|$ is a unit vector pointing to the closest point on the manifold. Therefore, for $\sigma_{d}=1$ we have:
$x^{\text{shift}}:=x+s_{d}(x)=x-\mathcal{D}_{\mathcal{M}}(x)\nabla_{x}\mathcal{D}_{\mathcal{M}}(x)/|\nabla_{x}\mathcal{D}_{\mathcal{M}}(x)|=G(F(x))=x^{*},$ (8)
@@
$s_{\text{non-u}}(x)=\nabla_{x}\log{P}_{\text{non-u}}\approx\nabla_{x}\log\hat{P}_{\text{non-u}}=:\hat{s}_{\text{non-u}}(x).$
@@
$\hat{s}_{\text{non-u}}(x)=-\frac{1}{2\sigma_{d}^{2}}\left(x-\bar{G}(x)\right),$ (9)
@@
Note that $\bar{G}(x)$, which is the mean of the contributions from all points in the manifold to the probability $P(x)$ does not necessarily lie on the manifold. In contrast, $x^{*}=G(F(x))$ is, by definition, a point on the manifold. See Fig. 3 for an illustration and Fig. 4 for a synthetic example. The integral over $z$ in the computation of $\bar{G}_{\alpha}(x)$ is approximated by randomly sampling the normal distribution centered around the training point $z_{\alpha}$ (see details in Appendix A).
A noisy or corrupted image $x$ can be viewed as a point in the ambient space. The image generation then becomes the task of finding an appropriate, though not necessarily orthogonal, projection of this point onto the manifold of clean, meaningful images. If the mappings and functions $G$, $F$, and $\mathcal{D}_{\mathcal{M}}$ are perfectly accurate, a single step can move $x$ closer to the corresponding clean image. Since the ambient space is sampled sparsely, especially in regions far from the manifold, the approximations of these mappings become less accurate as the distance from the manifold increases. To address this, we employ multiple iterative steps, gradually improving accuracy as we move closer to the manifold. This process resembles a diffusion-like flow; see Fig. 4 for an illustrative example. Equations 7 and 8 motivate a diffusion-like process guided by the distance function. The score defines a vector field in the ambient space. A step in the direction of the closest point on the manifold is
$x^{n+1}=x^{n}-\alpha\mathcal{D}_{\mathcal{M}}(x^{n})\nabla_{x}\mathcal{D}_{\mathcal{M}}(x^{n})/|\nabla_{x}\mathcal{D}_{\mathcal{M}}(x^{n})|\quad\text{\rm with}\quad 0<\alpha<1\quad\text{\rm and}\quad x^{0}={x}\ .$ (11)
Equation 11 does not take into account the distribution of training points on the manifold. To address this limitation, we combine it with the score of the kernel method to obtain:
$x^{n+1}=(1-\beta)x^{n}+\beta\bar{G}({x^{n}})-\alpha\mathcal{D}_{\mathcal{M}}(x^{n})\nabla_{x}\mathcal{D}_{\mathcal{M}}(x^{n})/|\nabla_{x}\mathcal{D}_{\mathcal{M}}(x^{n})|,$ (12)
where $0<\alpha,\beta$, $\alpha+\beta<1$, and $x^{0}={x}$. The trajectory of $x$ as it moves towards the manifold is illustrated in Fig. 4.
@@ 4.1 Manifold-Probabilistic Projection Model (MPPM)
$\mathcal{L}(F,G,\mathcal{D}_{\mathcal{M}})$ $=\lambda_{1}\sum_{x_{i}\notin\mathcal{M}}\left(\mathcal{D}_{\mathcal{M}}(x_{i})-\|x_{i}-x_{i}^{*}\|\right])^{2}+\lambda_{2}\sum_{x_{i}\in\mathcal{M}}\left(x_{i}^{\text{clean}}-G(F(x_{i}^{\text{clean}})\right)^{2}$ (13)
$+\lambda_{3}\sum_{x_{i}\in\mathcal{M}}|\mathcal{D}_{\mathcal{M}}(x_{i})|^{2}+\lambda_{4}\sum_{x_{i}\in\mathbb{R}^{D}}\left(\mathcal{D}_{\mathcal{M}}(x_{i})-|\mathcal{D}_{\mathcal{M}}(x_{i})|\right)^{2}+\lambda_{5}\sum_{x_{i}\in\mathbb{R}^{D}}\left(x_{i}^{\text{shift}}-x_{i}^{*}\right)^{2},$
where $x^{*}=G(F(x))$. The first term defines the distance function assuming a perfect autoencoder; the second is the standard autoencoder loss. The third term enforces the boundary condition on the distance function and the fourth ensures its positivity. The last term enforces the geometric consistency of Eq. (8) (see also Fig. 1).
@@ 4.2 Latent MPPM (LMPPM)
The key difference between the pixel space and the latent space is that, in the latter, we do not assume that encoded clean and meaningful images lie on a lower-dimensional manifold. Instead, we treat the set of encoded clean and meaningful images as a point cloud that occupies the full dimension of the latent space. We model this set as samples from a probability distribution $P(z)$. Let the set of clean and meaningful images be $\mathcal{X}^{\text{clean}}$ and the set of these encoded images be $S=\{F(\mathcal{X}^{\text{clean}})\}$. In this context, $S$ serves the role that the manifold $\mathcal{M}$ played in the previous section, in the sense that the distance function $\mathcal{D}_{S}$ is now computed in the latent space with respect to the set $S$. Let $x\in\mathbb{R}^{D}$ be an image and $z=F(x)\in\mathbb{R}^{d}$ its latent representation. The reconstructed image is then given by $\hat{x}=G(z)$. Let us define a distance function $\mathcal{\mathcal{D}_{S}}:\mathbb{R}^{d}\to\mathbb{R}$ such that $\mathcal{D}_{S}(z)$ measures the distance from $z$ to the set $S$ in the latent space. Using this, we define a shift in the latent space as: $z^{\text{shift}}:=z-\mathcal{\mathcal{D}_{S}}(z)\nabla_{z}\mathcal{\mathcal{D}_{S}}(z)/|\nabla_{z}\mathcal{\mathcal{D}_{S}}(z)|$. The loss function is then given by
$\mathcal{L}(F,G,\mathcal{D}_{S})$ $=\lambda_{1}\sum_{z_{i}\notin S}\left(\mathcal{D}_{S}(z_{i})-\|z_{i}-z_{i}^{*}\|\right])^{2}+\lambda_{2}\sum_{z_{i}\in S}\left(x_{i}^{\text{clean}}-G(z_{i})\right)^{2}$ (14)
@@
where $x_{i}^{*}=\arg\min_{\tilde{x}\in\mathcal{X}^{\text{clean}}}\|x_{i}-\tilde{x}\|$, and $z_{i}^{*}=F(x_{i}^{*})$. These definitions ensure that a generic point $x$ in the ambient space, whose closest clean image in the dataset is $x^{*}$ is mapped to $z=F(x)$ such that its nearest neighbor in $S$ is $z^{*}=F(x^{*})$. It is important to note that the set $S$ evolves over training iterations as the encoder $F$ and decoder $G$ are updated, and the distance function $\mathcal{D_{S}}$ is adjusted accordingly. The first three terms are the heart of the algorithm. The 4th element ensures positivity. The 5th and 6th terms improve consistency between all three networks. Ablation study empirically proves that these terms contribute to the performance of the method. By the kernel method, we obtain
@@
  $G,F,\mathcal{\mathcal{D}_{S}}\leftarrow\text{Train}\big(\mathcal{X}^{\text{clean}},\epsilon,\mathcal{L}(F,G,\mathcal{\mathcal{D}_{S}})\big)$   by 14
@@
   $z^{n+1}\leftarrow(1-\beta)z^{n}+\beta\bar{z}^{n}-\alpha\mathcal{\mathcal{D}_{S}}(z^{n})\nabla_{z}\mathcal{\mathcal{D}_{S}}(z^{n})/|\nabla_{z}\mathcal{\mathcal{D}_{S}}(z^{n})|$   by 15
@@ 5 Experiments
We evaluated our MPPM method on synthetic data and our LMPPM method on real-world image datasets, where we simultaneously trained an autoencoder-like network for $F$ and $G$, and a different network for the distance function $\mathcal{D}_{\mathcal{M}}$ and $\mathcal{D}_{S}$. It is important to note that training was performed exclusively with Gaussian noise degradation, while at inference time we evaluated the models under a variety of other degradation types. We compared our results with standard denoising autoencoders (DAE) (denoising_original_vincent2008extracting, ) and latent diffusion models (LDM) (rombach2022ldm, ). For synthetic experiments, we evaluated on a one-dimensional manifold: a half-circle lying in the xy plane and embedded in $\mathbb{R}^{3}$. The points in the circle are sampled according to angular coordinates drawn from truncated normal distributions (see Figs. 4 and 5).
For real-world data, we experiment with MNIST (lecun1998mnist, ) and the SCUT-FBP5500 facial beauty dataset (liang2017SCUT, ). To evaluate restoration performance, we apply three types of degradation to MNIST: Gaussian noise, downsampling (super-resolution), and elastic deformation, each at two severity levels. For SCUT-FBP5500, we consider four types of degradation: Gaussian noise, downsampling, random scribbles, and black patches (inpainting), also applied at two severity levels.
@@
We trained our proposed methods and the comparison baselines to assess their performance in the different datasets. Detailed architecture specifications and hyperparameters are provided in Appendices B and C. For synthetic data, we implement MPPM using MLP architectures. For MNIST, we employ a CNN-based autoencoder for both DAE and our LMPPM method, while for SCUT-FBP5500 we adopt a U-Net architecture with skip connections. In addition, we construct an extra set of skip connections from the latent space and combine them with the original skips through weighted summation (see Appendix B). The distance functions $\mathcal{D}_{\mathcal{M}}$ and $\mathcal{D}_{S}$ are implemented as MLPs with progressively decreasing layer sizes to perform dimensionality reduction. For LDM, we integrate the corresponding DAE backbone (in place of the autoencoder) with a standard diffusion model, using 2000 diffusion steps.

### v2

@@ 3 Geometric View of Diffusion Models
Since both the encoder $F$ and the manifold distance $\mathcal{D}_{\mathcal{M}}$ are defined over the ambient space $\mathbb{R}^{D}$, training mappings that induce a diffusion-like flow from corrupted images back to clean manifold samples requires sampling in this high-dimensional space. However, this sampling is inherently difficult due to the curse of dimensionality. Following diffusion models, we obtain ambient samples by adding Gaussian noise to the data points. Although this strategy does not cover all possible corruptions, it empirically yields effective mappings. Notably, despite being trained solely with Gaussian noise, our models generalize well to other types of image corruptions at test time.
@@
To relate a corrupted image to its clean projection, we use the concept of the score. The score is a $D$-dimensional vector field defined by $s(x)=\nabla_{x}\log P(x)$, which points in the direction of the steepest ascent of the probability density. For the distance-based probability distribution $P_{d}(x)$ (6), we obtain:
@@
Since $\mathcal{D}_{\mathcal{M}}(x)$ is the distance to the manifold, its gradient $\nabla_{x}\mathcal{D}_{\mathcal{M}}(x)$ is a unit vector pointing to the closest point on the manifold. Therefore, for $\sigma_{d}=1$ we have:
$x^{\text{shift}}:=x+s_{d}(x)=x-\mathcal{D}_{\mathcal{M}}(x)\nabla_{x}\mathcal{D}_{\mathcal{M}}(x)=G(F(x))=x^{*},$ (10)
@@
$s(x)=\nabla_{x}\log{P}{{}_{\text{non-u}}(x)}\approx\nabla_{x}\log\hat{P}{{}_{\text{non-u}}(x)}=:\hat{s}(x).$
@@
$\hat{s}(x)=-\frac{1}{2\sigma_{d}^{2}}\left(x-\bar{G}(x)\right),$ (11)
@@
$\bar{G}_{\alpha}(x)={\frac{1}{\hat{P}_{\text{non-u}}(x)Q_{\text{ker}}}}\int\left[G(z)P(x\mid G(z))\exp\left({-\frac{\|z-z_{\alpha}\|^{2}}{2\sigma_{\text{ker}}^{2}}}\right)\right]dz.$ (12)
@@
Compared to prior work, $\bar{G}(x)$ in Eq. (11) is the same object as $\mathbb{E}_{y\in P_{c}}[y|x]$ in Kadkhodaie et al (Kadkhodaie et al., 2023). This quantity is intractable because the integral can’t be approximated by sampling clean images beyond the points in the data set. It is therefore often replaced by the denoiser $x^{*}$ (Kadkhodaie et al., 2023). Clearly, the approximation of $\bar{G}$ by $x^{*}$ is justified only under a uniform distribution over the manifold. In contrast, in our geometric formulation the integral over $z$ in the computation of $\bar{G}_{\alpha}(x)$ is approximated by randomly sampling the normal distribution centered around the training point $z_{\alpha}$ (for details, see Appendix A.2.).
A noisy or corrupted image $x$ can be viewed as a point in the ambient space. The image generation then becomes the task of finding an appropriate, though not necessarily orthogonal, projection of this point onto the manifold of clean, meaningful images. If the mappings and functions $G$, $F$, and $\mathcal{D}_{\mathcal{M}}$ are perfectly accurate, a single step can move $x$ closer to the corresponding clean image. Since the ambient space is sampled sparsely, particularly in regions far from the manifold, the learned approximations become less accurate as the distance from the manifold increases. To mitigate this, we apply multiple iterative steps that progressively refine the estimate as the trajectory approaches the manifold. This procedure resembles a diffusion-like flow; see Fig. 4 for an illustrative example.
Equations (9) and (10) motivate a diffusion-like process guided by the distance function. The score induces a vector field in the ambient space. Using the Tweedie formula (Efron, 2011), we update the sample by taking a step toward the manifold projection, i.e., toward the closest point on the manifold. This yields the iterative update
$x^{n+1}=x^{n}-\alpha\mathcal{D}_{\mathcal{M}}(x^{n})\hat{\nabla}_{x}\mathcal{D}_{\mathcal{M}}(x^{n})\quad\text{\rm with}\quad 0<\alpha<1\quad\text{\rm and}\quad x^{0}={x}$ (13)
where
$\hat{\nabla}_{x}\mathcal{D}_{\mathcal{M}}(x):=\frac{\nabla_{x}\mathcal{D}_{\mathcal{M}}(x)}{\|\nabla_{x}\mathcal{D}_{\mathcal{M}}(x)\|}$ (14)
denotes the normalized gradient of the distance function. Since the distance network provides only an approximation, we normalize the gradient to better control the step size.
Equation (13) does not take into account the distribution of training samples along the manifold. To address this limitation, we incorporate the score induced by the kernel formulation and, by applying Tweedie formula (see Appendix A.3), to obtain
$x^{n+1}=(1-\beta)x^{n}+\beta\bar{G}({x^{n}})-\alpha\mathcal{D}_{\mathcal{M}}(x^{n})\hat{\nabla}x\mathcal{D}\mathcal{M}(x^{n}),\quad\text{with }0<\alpha,\beta,\ \alpha+\beta<1,\ \ x^{0}=x.$ (15)
The trajectory of $x$ as it moves toward the manifold is illustrated in Fig. 4.
@@ 4.1 Manifold-Probabilistic Projection Model (MPPM)
$\mathcal{L}(F,G,\mathcal{D}_{\mathcal{M}})$ $=\lambda_{1}\sum_{x_{i}\notin\mathcal{M}}\left(\mathcal{D}_{\mathcal{M}}(x_{i})-\|x_{i}-x_{i}^{*}\|\right])^{2}+\lambda_{2}\sum_{x_{i}\in\mathcal{M}}\left\|x_{i}^{\text{clean}}-G(F(x_{i}^{\text{clean}}))\right\|^{2}$ (16)
$+\lambda_{3}\sum_{x_{i}\in\mathcal{M}}|\mathcal{D}_{\mathcal{M}}(x_{i})|^{2}+\lambda_{4}\sum_{x_{i}\in\mathbb{R}^{D}}\left(\mathcal{D}_{\mathcal{M}}(x_{i})-|\mathcal{D}_{\mathcal{M}}(x_{i})|\right)^{2}+\lambda_{5}\sum_{x_{i}\in\mathbb{R}^{D}}\rVert x_{i}^{\text{shift}}-x_{i}^{*}\rVert^{2},$
where
$x_{i}^{\text{shift}}=x_{i}-\mathcal{D}_{\mathcal{M}}(x_{i})\hat{\nabla}_{x}\mathcal{D}_{\mathcal{M}}(x_{i})$ (17)
and
$x_{i}^{*}=G(F(x_{i})).$ (18)
The first term defines the distance function assuming a perfect autoencoder; the second is the standard autoencoder loss. The third term enforces the boundary condition on the distance function and the fourth ensures its positivity. The last term enforces the geometric consistency of Eq. (10) (see also Fig. 1). At inference, we use the Tweedie formula (15), where
$\bar{G}_{\alpha}(x)\approx\frac{\sum_{z_{i}\in\mathcal{N}(z_{\alpha},\sigma_{\text{ker}}^{2})}G(z_{i})\exp\left(-{\|{x}-G(z_{i})\|^{2}/2\sigma_{d}^{2}}\right)}{\sum_{z_{i}\in\mathcal{N}(z_{\alpha},\sigma_{\text{ker}}^{2})}\exp\left({-\|{x}-G(z_{i})\|^{2}/2\sigma_{d}^{2}}\right)}$ (19)
as derived in Appendix A.2.
@@
Figure 6: Left: Diffusion model trajectory, Right: MPPM trajectory
Figure 7: Left: Diffusion restoration, MSE=0.138, Right: MPPM restoration, MSE=0.026
Figures 6 and 7 illustrate the advantages of the proposed MPPM method compared to the diffusion model. We used 1000 diffusion steps during both training and inference. Because the data points are not uniformly distributed, most of the diffusion model’s reconstructed samples concentrate in the dense region (the upper half-circle), as shown in the left panel of Fig. 7. In contrast, the proposed MPPM method effectively handles this non-uniformity through our formulation, resulting in a significantly smaller reconstruction error. Pseudocode for DDPM (Denoising Diffusion Probabilistic Model) training and inference is summarized in Algorithms 2 and 3. Note the difference from Chen et.al. 2024 (Chen et al., 2024), who analyzed the trajectory of the flow by measuring the deviation from the line between the degraded image and the point found on the manifold.
Algorithm 2 DDPM Training
1:Precompute noise schedule:
[…]
5:for each batch do
6:  Sample timestep $t\sim\text{Uniform}(0,T-1)$
7:  Sample noise $\epsilon\sim\mathcal{N}(0,I)$
8:  Forward diffusion:
$x_{t}=\sqrt{\bar{\alpha}_{t}}\,x_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\epsilon$
9:  Predict noise: $\epsilon_{\text{pred}}=\text{model}(x_{t},\,t/T)$
10:  Compute loss:
$L=\|\epsilon-\epsilon_{\text{pred}}\|^{2}$
11:  Backpropagate and update parameters
12:end for
Algorithm 3 DDPM Inference (Stochastic, 1000 steps)
1:Initialize $x_{T}\sim\mathcal{N}(0,I)$
2:for t=T1t=T-1 down to 0 do
3:  Predict noise: $\epsilon_{\text{pred}}=\text{model}(x_{t},\,t/T)$
4:  Denoise:
$x_{t}=\frac{x_{t}-\text{coef}_{t}\,\epsilon_{\text{pred}}}{\sqrt{\alpha_{t}}}$
5:  if $t>0$ then
6:   Sample $z\sim\mathcal{N}(0,I)$
7:   Add noise: $x_{t}=x_{t}+\sigma_{t}z$
8:  end if
9:end for
10:return $x_{0}$
@@ 4.2 Latent MPPM (LMPPM)
The key difference between the pixel space and the latent space is that, in the latter, we do not assume that encoded clean and meaningful images lie on a lower-dimensional manifold. Instead, we treat the set of encoded clean and meaningful images as a point cloud that occupies the full dimension of the latent space. We model this set as samples from a probability distribution $P(z)$. Let the set of clean and meaningful images be $\mathcal{X}^{\text{clean}}$ and the set of these encoded images be $S=\{F(\mathcal{X}^{\text{clean}})\}$. In this context, $S$ serves the role that the manifold $\mathcal{M}$ played in the previous section, in the sense that the distance function $\mathcal{D}_{S}$ is now computed in the latent space with respect to the set $S$. Let $x\in\mathbb{R}^{D}$ be an image and $z=F(x)\in\mathbb{R}^{d}$ its latent representation. The reconstructed image is then given by $\hat{x}=G(z)$. Let us define a distance function $\mathcal{\mathcal{D}_{S}}:\mathbb{R}^{d}\to\mathbb{R}$ such that $\mathcal{D}_{S}(z)$ measures the distance from $z$ to the set $S$ in the latent space. Using this, we define a shift in the latent space as: $z^{\text{shift}}:=z-\mathcal{\mathcal{D}_{S}}(z)\hat{\nabla}_{z}\mathcal{\mathcal{D}_{S}}(z)$. The loss function is then given by
$\mathcal{L}(F,G,\mathcal{D}_{S})$ $=\lambda_{1}\sum_{z_{i}\notin S}\left(\mathcal{D}_{S}(z_{i})-\|z_{i}-z_{i}^{*}\|\right])^{2}+\lambda_{2}\sum_{z_{i}\in S}\left\|x_{i}^{\text{clean}}-G(z_{i})\right\|^{2}$ (20)
@@
where $x_{i}^{*}=\arg\min_{\tilde{x}\in\mathcal{X}^{\text{clean}}}\|x_{i}-\tilde{x}\|$, and $z_{i}^{*}=F(x_{i}^{*})$. These definitions ensure that a generic point $x$ in the ambient space, whose closest clean image in the dataset is $x^{*}$ is mapped to $z=F(x)$ such that its nearest neighbor in $S$ is $z^{*}=F(x^{*})$. It is important to note that the set $S$ evolves over training iterations as the encoder $F$ and decoder $G$ are updated, and the distance function $\mathcal{D_{S}}$ is adjusted accordingly. The first three terms are the heart of the algorithm. The 4th element ensures positivity. The 5th and 6th terms improve consistency between all three networks. Ablation study empirically proves that these terms contribute to the performance of the method. By the kernel method, $\bar{G}(x)$ is replaced by
@@
  $G,F,\mathcal{\mathcal{D}_{S}}\leftarrow\text{Train}\big(\mathcal{X}^{\text{clean}},\epsilon,\mathcal{L}(F,G,\mathcal{\mathcal{D}_{S}})\big)$   by (20)
@@
   $z^{n+1}\leftarrow(1-\beta)z^{n}+\beta\bar{z}^{n}-\alpha\mathcal{\mathcal{D}_{S}}(z^{n})\hat{\nabla}_{z}\mathcal{\mathcal{D}_{S}}(z^{n})$   by (21)
@@ 5 Experiments
We evaluated our MPPM method on synthetic data and our LMPPM method on real-world image datasets, where we simultaneously trained an autoencoder-like network for $F$ and $G$, and a different network for the distance function $\mathcal{D}_{\mathcal{M}}$ and $\mathcal{D}_{S}$. It is important to note that training was performed exclusively with Gaussian noise degradation, while at inference time we evaluated the models under a variety of other degradation types. We compared our results with standard denoising autoencoders (DAE) (Vincent et al., 2008) and latent diffusion models (LDM) (Rombach et al., 2022). For synthetic experiments, we evaluated on a one-dimensional manifold: a half-circle lying in the xy plane and embedded in $\mathbb{R}^{3}$. The points in the circle are sampled according to angular coordinates drawn from truncated normal distributions (see Figs. 4 and 5).
For real-world data, we experiment with MNIST (LeCun, 1998), SCUT-FBP5500 facial beauty dataset Liang et al. (2018), and CelebA-HQ-256Liu et al. (2015). To evaluate restoration performance, we apply three types of degradation to MNIST: Gaussian noise, downsampling (super-resolution), and elastic deformation, each at two severity levels. For SCUT-FBP5500 and CelebA-HQ-256, we consider four types of degradation: Gaussian noise, downsampling, random scribbles, and black patches (inpainting), also applied at two severity levels.
@@
We trained our proposed methods and the comparison baselines to assess their performance in the different datasets. Detailed architecture specifications and hyperparameters are provided in appendices B and C. For synthetic data, we implement MPPM using MLP architectures. For MNIST, we employ a CNN-based autoencoder for both DAE and our LMPPM method, while for SCUT-FBP5500 and CelebA-HQ-256 we adopt a U-Net architecture with skip connections. In addition, we construct an extra set of skip connections from the latent space and combine them with the original skips through weighted summation (see Appendix B). The distance functions $\mathcal{D}_{\mathcal{M}}$ and $\mathcal{D}_{S}$ are implemented as MLPs with progressively decreasing layer sizes to perform dimensionality reduction. For LDM, we integrate the corresponding DAE backbone (in place of the autoencoder) with a standard diffusion model, using 2000 diffusion steps.